### v1

@@ 2.2. Roe algebras and their ideals
The algebra Roe algebra of $X$, denoted by $\mathbb{C}[X,\mathcal{H}_{X}]$ (or $\mathbb{C}[X]$ for simplicity), is a subalgebra of $\mathcal{B}(\mathcal{H}_{X})$ consisting of all operators $T$ satisfying:
@@
We should remark that the definition of the Roe algebra does not depend on the choice of $\mathcal{H}$ up to a non-canonical isomorphism. The following lemma is proved in [STY2002, Lemma 4.4], which reveals that Roe algebras can be fundamentally understood as groupoid $C^{*}$-algebras. If we replace $\mathcal{H}$ with $\mathbb{C}$, the resulting Roe algebra is called the uniform Roe algebra, denoted by $C^{*}_{u}(X)$.
@@
Since $\mathcal{I}$ is an ideal, one can prove that $U(\mathcal{I})$ is an invariant open subset, see [CW2004] for a detailed discussion.
@@
  • After fixing the support, the coefficient part of each operator (i.e., the $\ell^{\infty}(X,\mathcal{K})$ component) further determines the rank distribution of operator entries within the ideal, as discussed in [CW2006, WFZ2025].
In this paper, we will not consider the $\ell^{\infty}(X,\mathcal{K})$ aspect. That is, we assume the ideal $\mathcal{I}$ has maximal rank distribution over its corresponding invariant open set, i.e., $\mathcal{I}$ contains all rank distributions supported on $U(\mathcal{I})$. We will not review the precise definition of rank distribution here; interested readers may refer to [CW2006, WFZ2025] for details.
@@
The ghostly ideal of $C^{*}_{u}(X)$ with respect to $U$, denoted by $\mathcal{I}_{\mathcal{G}}(X,U)$, is defined to be
@@
One should notice that $\mathbb{I}[X,U]$ is dense in $\mathcal{I}(X,U)$ and is also an ideal of $\mathbb{C}[X]$. Note that the definitions of geometric ideals and ghostly ideals are very similar, but they have essential differences. The geometric ideal is constructed in the algebraic Roe algebra by taking the part where the $\varepsilon$-support set lies in $U$ and then generating an ideal, while the ghostly ideal is obtained in the completed Roe algebra by taking the part where the $\varepsilon$-support set lies in $U$. The essential difference lies in the order of taking the completion and imposing the support condition. Therefore, it is straightforward to see that the geometric ideal is a subalgebra of the ghostly ideal. Moreover, the ghostly ideal is larger than the corresponding geometric ideal, and the two ideals coincide only when the space has Yu’s Property A (see [WZ2023]).
@@
Let $X$ be a metric space with bounded geometry. If $X$ admits a coarse embedding into some $\ell^{p}$-space with $p\in[1,\infty)$, then for any invariant open subset $U$, the inclusion $i:\mathcal{I}(X,U)\to\mathcal{G}(X,U)$ induces an isomorphism on $K$-theory, i.e.,
$i_{*}:K_{*}(\mathcal{I}(X,U))\to K_{*}(\mathcal{G}(X,U))$
@@
As a result of P. Nowak [Nowak2006], for $p\in[1,2)$, coarse embeddability into $\ell^{p}$ is equivalent to coarse embeddability into Hilbert space. In this case, the result is proved in [WFZ2025]. Therefore, we may assume $p>2$ and that there exists a coarse embedding $f:X\to\ell^{p}$. Without loss of generality, we shall assume $p\geq 2$ throughout this paper.
It is worth mentioning that for $q\in[2,\infty), p\ell^{p}$ admits a coarse embedding into $\ell^{q}$ if and only if $p\leq q$, see [MN2008, Theorem 1.10]. Thus, coarse embedding into $\ell^{p}$-space is stronger than coarse embedding into Hilbert space when $p>2$. In other words, for any metric space $X$ with bounded geometry, there exists $p_{X}\in[2,\infty]$ such that $X$ admits a coarse embedding into $\ell^{p}$ for any $p\geq p_{X}$. There is still no known example for a space with bounded geometry that coarsely embeds into $\ell^{p}$ for $p>2$ but can not coarsely embed into Hilbert space.
@@ 2.3. Coarse invariant of ghostly/geometric ideals
In this subsection, we shall show that the ghostly ideal is coarse invariant.
@@
If $f:X\to Y$ is a coarse equivalent between the pair $(X,U)$ and $(Y,V)$, then $f$ induces an isometry $W_{f}:\mathcal{H}_{X}\to H_{Y}$ such that
@@
By [HIT2020, Proposition 4.3.4 & Proposition 4.3.5], one can find a covering isometry $W_{f}:\mathcal{H}_{X}\to\mathcal{H}_{Y}$ for $f$. For discrete metric spaces with bounded geometry, the construction of $W_{f}$ will be easier. We shall briefly recall the process here.
@@
Conjugation by $U$ induces the identity map on $K$-theory and one can easily chech that $U\alpha_{1}(T)U^{*}=\alpha_{2}(T)$. This completes the proof.
@@ 2.4. Reduction to sparse spaces
By [CWY2013, Section 4], one can find an $\omega$-excisive cover $\{Y,Z\}$ of $X$ such that $Y$, $Z$, and $Y\cap Z$ are all sparse, i.e., they are all coarse disjoint unions of finite spaces. From the definition of Stone-Čech compactification, we conclude that $\beta X=\beta Y\cup\beta Z$. Then the canonical inclusion $\iota:Y\to X$ induces a canonical inclusion $\iota:G(Y)\to G(X)$. Actually, one can easily check the identification $G(Y)=G(X)|^{\beta Y}_{\beta Y}$.
@@
Write $V_{Y}$ the $G(X)$-invariant open subset in $\beta X$. Notice that $\mathcal{I}(X,V_{Y})$ is isomorphic to the ideal of $C^{*}(X)$ generated by $Y$, denoted by $C^{*}(X,Y)$ as in [HRY1993]. Thus, one has that
@@
With notation as above, we have that the canonical inclusion map $\iota:C^{*}(Y)\cap\mathcal{I}\to\mathcal{I}(X,V_{Y})\cap\mathcal{I}$ induces an isomorphism on $K$-theory
@@
By 2.8, it induces an isomorphism on $K$-theory, i.e.,
@@
If 2.6 holds for all sparse subspaces $Y$ of $X$, then it holds for $X$ itself.
@@
Take $\{Y,Z\}$ to be an $\omega$-excisive cover of $X$ such that $Y,Z$ and $Y\cap Z$ are all sparse. Since $Y\cup Z=X$, we have $V_{Y}\cup V_{Z}=\beta X$ by the definition of the Stone-Čech compactification, where $V_{Y}$ and $V_{Z}$ are $G(X)$-invariant open subset generated by $Y$ and $Z$, respectively. It is proved in [HRY1993] that
@@
induces an isomorphism on $K$-theory for $W=Y,Z$ and $Y\cap Z$ which are all sparse. By 2.10, we conclude that $i_{*}$ is identified with
@@
By 2.9, the map $i_{*}$ is further identified with
@@
Therefore, it suffices to prove 2.6 for the case when $X$ is a sparse space, as the general case will then follow. This completes the proof of the proposition. ∎
@@ 3. The $\ell^{p}$-Bott-Dirac construction and twisted ghostly ideals
In the following subsections, we shall prove 2.6 in several steps. First of all, we shall briefly recall the definition of the Bott-Dirac operator in finite-dimensional $\ell^{p}$-spaces. Please refer to [WXYZ2024] for a detailed introduction.
@@
When $p\geq 2$, the extended Mazur map is Lipschitz with constant $L$.
@@
For any $v\in(E,\|,\|_{p})$, we define the Bott generator to be the unbounded function
@@
We summarize the facts we will need about the Bott-Drac operator in the next result. A detailed proof for the proposition can be founded in [WXYZ2024, Lemma 2.7]:
Proposition 3.1 ([WXYZ2024]).
@@
The following bounded version of the Bott-Dirac operator will be frequently used in what follows.
@@
Notice that $F_{s,v}$ is a bounded operator on $L^{2}(E,\mathfrak{h}_{E})$. There is a natural action of $C_{0}(E)$ on $L^{2}(E,\mathfrak{h}_{E})$, given by pointwise multiplication. Consequently, for any $p\in[1,\infty)$, we can discuss the propagation of Fs,vF{s,v} in the $E$-direction with respect to the norm $|\cdot|_{p}$. We shall denote this number by $\textup{Prop}_{E,p}(F_{s,v})$.
For any $v\in E$ and $R>0$, we denote $\chi_{v,R}:=\chi_{B(v,R)}$ for simplicity. The following theorem summarizes some properties of $F_{s,v}$. For the detailed proof, we refer the reader to [WXYZ2024, Proposition 2.17].
Proposition 3.3.
For each $\varepsilon>0$, there exists an odd function $\Phi:\mathbb{R}\to[-1,1]$ such that $\Phi(t)\to\pm 1$ as $t\to\pm\infty$ with the following properties.
@@
    For any $s\in[1,\infty)$ and all $v\in E$, $\|F_{s,v}-\Phi(B_{s,v})\|<\varepsilon$.
@@
    There exists $R_{0}>0$ such that for all $s\in[1,\infty)$, $p\geq 2$ and $v\in E$, $\textup{Prop}_{E,p}(\Phi(B_{s,v}))\leq s^{-1}R_{0}$.
@@
    For any $s\in[1,\infty)$ and all $v\in E$, the operator $\Phi(B_{s,v})^{2}-1$ is compact.
@@
    For any $s\in[1,\infty)$ and all $v,w\in E$, the operator $\Phi(B_{s,v})-\Phi(B_{s,w})$ is compact.
  5. (5)
    There exists $c>0$ such that for all $s\in[1,\infty)$, $p\geq 2$ and all $v,w\in E$, Φ(B_s,v)-Φ(B_s,w)∥≤cv-w_p.
  6. (6)
    For any $v\in E$, the function [1,)B(L^2(E,h_E))  sΦ(B_s,v) is continuous under the strong operator topology.
  7. (7)
    The family of functions [1,)B(L^2(E,h_E))  sΦ(B_s,v)^2-1 is norm equi-continuous as $v$ varies over $E$ and $s$ varies over any fixed compact subset of $1,\infty)$.
  8. (8)
    For any $r\geq 0$, the family of functions [1,)B(L^2(E,h_E))  sΦ(B_s,v)-Φ(B_s,w) is norm equi-continuous as $(v,w)$ varies over $E\times E$ with $|v-w|\leq r$, and $s$ varies over any fixed compact subset of $1,\infty)$.
  9. (9)
    There exists $R_{1}>0$ such that for all $R\geq R_{1}$, $s\in[L\cdot(\operatorname{dim}E_{n})^{2},\infty)$, and $v\in E$ we have that: ——(Φ(B_s,v)^2-1)(1-χ_v,R)——¡3ε.
  10. (10)
    For any $r\geq 0$, there exists $R_{2}>0$ such that for all $R\geq R_{2}$, $s\in[L\cdot(\operatorname{dim}E_{n})^{2},\infty)$, and $v,w\in E\times E$ with $|x-y|\leq r$ we have that: ——(Φ(B_s,v)-Φ(B_s,w))(1-χ_v,R)——¡3ε.
We remark that the choice of $R_{1}$ in (9) and the choice of $R_{2}$ in (10) do not depend on the dimension of EE.∎
@@
The first step ensures that, without loss of generality, we may assume $X=\bigsqcup_{n=0}^{\infty}X_{n}$ is a sparse space that admits a coarse embedding into $\ell^{p}$. By [HIT2020, Lemma 12.5.4], we can assume the distance between $X_{n}$ and $X_{m}$ is infinite whenever $n\neq m$ (called separated disjoint union of $\{X_{n}\}_{n\in\mathbb{N}}$). In this case, the Hilbert space $\mathcal{H}_{X}$ has a canonical decomposition $\mathcal{H}_{X}=\bigoplus_{n\in\mathbb{N}}\mathcal{H}_{X_{n}}$. The Roe algebra $C^{*}(X)$ is therefore a subalgebra of $\prod_{n\in\mathbb{N}}C^{*}(X_{n})$
@@
Next, we will use the bounded Bott-Dirac map from 3.2 to construct the index map:
@@
For each $s\in[1,\infty)$, $F_{s,v,E_{n}}:L^{2}(E_{n},\mathfrak{h}_{E_{n}})\to L^{2}(E_{n},\mathfrak{h}_{E_{n}})$ is the bounded operator as in 3.2. Set $s_{n}:=L(\operatorname{dim}E_{n})^{2}$, where $L$ is the constant in 3.3.
@@
Analogously, if $\Phi$ has the properties in 3.3 for some $\varepsilon>0$, we write $F^{\Phi}$, $F^{\Phi}_{s}$ and $F^{\Phi}_{s,n}$ for the operators built in the same way as $F$, $F_{s}$ and $F_{s,n}$, but replacing $F_{s,v,E_{n}}$ with $\Phi(B_{s,v,E_{n}})$.
By 3.3, the function $F$ is strong *-continuous with respect to $s\in[1,\infty)$. To introduce the index map, we also need the following lemma. The proof of the lemma can be found in [WXYZ2024, Proposition 2.19] and [HIT2020, Section 12.3]. We will omit some details here.
@@
It should be noted that in 3.5, the parameter $s$ is shifted to the right by $s_{n}$ (depending on the dimension) to ensure the validity of the preceding 3.6 applying (9), (10) in 3.3.
By [HIT2020, Corollary 2.7.4], a $K_{0}$-theory element of $C^{*}(X)$ can be represented by a single projection $[P]$. For any projection $P\in C^{*}(\mathcal{H}_{X})$, $PF_{s}P$ is an odd, self-adjoint operator on the graded Hilbert space $L^{2}[1,\infty)\otimes\mathcal{H}_{X,E}$. We can build an index class of $K_{0}(PA(X,E)P)$ by using $PF_{s}P$ and 3.6. We can decompose $PF_{s}P$ as a matrix with respect to the grading on each $\mathcal{H}_{X,E}$ and $PF_{s}P$ has the form
@@
The index map is a well-defined group homomorphism. For each $s\in[1,\infty)$, define $\iota^{s}:A(X,E)\to C^{*}(X,H_{X,E})$ to be the evaluation map $(T_{s})\mapsto T_{s}$. The following proposition is proved in [WXYZ2024, Proposition 2.20], also see [HIT2020, Section 12.3]. We shall provide a detailed proof in 3.12.
@@
We will only prove the case for the map $s\mapsto T(F_{s}^{2}-1)$ by using (7) in 3.3, as the argument for $T,F_{s}]$ is analogous by using (8) in 3.3. By 3.6, we already have these two functions belong to $A(X,E)$.
@@
Combining this with 3.10, we have that

### v2

@@ 2.2. Roe algebras and their ideals
The algebraic Roe algebra of $X$, denoted by $\mathbb{C}[X,\mathcal{H}_{X}]$ (or $\mathbb{C}[X]$ for simplicity), is a subalgebra of $\mathcal{B}(\mathcal{H}_{X})$ consisting of all operators $T$ satisfying:
@@
We should remark that the definition of the Roe algebra does not depend on the choice of $\mathcal{H}$ up to a non-canonical isomorphism. The following lemma is proved in [Skandalis et al., 2002, Lemma 4.4], which reveals that Roe algebras can be fundamentally understood as groupoid $C^{*}$-algebras. If we replace $\mathcal{H}$ with $\mathbb{C}$, the resulting Roe algebra is called the uniform Roe algebra, denoted by $C^{*}_{u}(X)$.
@@
Since $\mathcal{I}$ is an ideal, one can prove that $U(\mathcal{I})$ is an invariant open subset, see Chen and Wang [2004] for a detailed discussion.
@@
  • After fixing the support, the coefficient part of each operator (i.e., the $\ell^{\infty}(X,\mathcal{K})$ component) further determines the rank distribution of operator entries within the ideal, as discussed in Chen and Wang [2006]; Wang et al. [2025].
In this paper, we will not consider the $\ell^{\infty}(X,\mathcal{K})$ aspect. That is, we assume the ideal $\mathcal{I}$ has maximal rank distribution over its corresponding invariant open set, i.e., $\mathcal{I}$ contains all rank distributions supported on $U(\mathcal{I})$. We will not review the precise definition of rank distribution here; interested readers may refer to Chen and Wang [2006]; Wang et al. [2025] for details.
@@
The ghostly ideal of $C^{*}(X)$ with respect to $U$, denoted by $\mathcal{G}(X,U)$, is defined to be
@@
One should notice that $\mathbb{I}[X,U]$ is dense in $\mathcal{I}(X,U)$ and is also an ideal of $\mathbb{C}[X]$. Note that the definitions of geometric ideals and ghostly ideals are very similar, but they have essential differences. The geometric ideal is constructed in the algebraic Roe algebra by taking the part where the $\varepsilon$-support set lies in $U$ and then generating an ideal, while the ghostly ideal is obtained in the completed Roe algebra by taking the part where the $\varepsilon$-support set lies in $U$. The essential difference lies in the order of taking the completion and imposing the support condition. Therefore, it is straightforward to see that the geometric ideal is a subalgebra of the ghostly ideal. Moreover, the ghostly ideal is larger than the corresponding geometric ideal, and the two ideals coincide only when the space has Yu’s Property A (see Wang and Zhang [2025]).
@@
Let $X$ be a metric space with bounded geometry. If $X$ admits a coarse embedding into some $\ell^{p}$-space with $p\in[1,\infty)$, then for any invariant open subset $U\subseteq\beta X$, the inclusion $i:\mathcal{I}(X,U)\to\mathcal{G}(X,U)$ induces an isomorphism on $K$-theory, i.e.,
$i_{*}:K_{*}(\mathcal{I}(X,U))\to K_{*}(\mathcal{G}(X,U))$ (1)
@@
Remark 2.7.
By the following six-term exact sequence
$K_{0}(\mathcal{I}(X,U))$$K_{0}(\mathcal{G}(X,U))$$K_{0}\left(\frac{\mathcal{G}(X,U)}{\mathcal{I}(X,U)}\right)$$K_{1}\left(\frac{\mathcal{G}(X,U)}{\mathcal{I}(X,U)}\right)$$K_{1}(\mathcal{G}(X,U))$$K_{1}(\mathcal{I}(X,U)),$}
we conclude that the map $i_{*}$ in (1) is an isomorphism if and only if $K_{*}(\frac{\mathcal{G}(X,U)}{\mathcal{I}(X,U)})=0$. As a result, for an $\ell^{p}$-coarsely embeddable space $X$, one has that $K_{*}(\frac{\mathcal{G}(X,U)}{\mathcal{I}(X,U)})=0$ for any invariant open subset $U\subseteq\beta X$.
As a result of P. Nowak Nowak [2006], for $p\in[1,2]$, coarse embeddability into $\ell^{p}$ is equivalent to coarse embeddability into Hilbert space. In this case, the theorem is proved in Wang et al. [2025]. It is worth mentioning that for $q\in[2,\infty), p\ell^{p}$ admits a coarse embedding into $\ell^{q}$ if and only if $p\leq q$, see [Mendel and Naor, 2008, Theorem 1.10]. Thus, coarse embedding into $\ell^{p}$-space is weaker than coarse embedding into Hilbert space when $p>2$. However, $\ell^{p}$ is not of bounded geometry and there is still no known example of a metric space with bounded geometry that coarsely embeds into $\ell^{p}$ for $p>2$ but can not coarsely embed into Hilbert space.
@@ 2.3. Coarse invariant of ghostly/geometric ideals
In this subsection, we shall show that the ghostly ideal is coarsely invariant. The reader is referred to [Jiang and Zhang, 2025, Section 6] for some relevant discussion.
@@
If $f:X\to Y$ is a coarse equivalence between the pair $(X,U)$ and $(Y,V)$, then $f$ induces an isometry $W_{f}:\mathcal{H}_{X}\to H_{Y}$ such that
@@
By [Willett and Yu, 2020, Proposition 4.3.4 & Proposition 4.3.5], one can find a covering isometry $W_{f}:\mathcal{H}_{X}\to\mathcal{H}_{Y}$ for $f$. For discrete metric spaces with bounded geometry, the construction of $W_{f}$ will be easier. We shall briefly recall the process here.
@@
Conjugation by $U$ induces the identity map on $K$-theory, and one can easily check that $U\alpha_{1}(T)U^{*}=\alpha_{2}(T)$. This completes the proof.
@@ 2.4. Reduction to sparse spaces
By [Chen et al., 2013, Section 4], one can find an $\omega$-excisive cover $\{Y,Z\}$ of $X$ such that $Y$, $Z$, and $Y\cap Z$ are all sparse, i.e., they are all coarse disjoint unions of finite spaces. From the definition of Stone-Čech compactification, we conclude that $\beta X=\beta Y\cup\beta Z$. Then the canonical inclusion $\iota:Y\to X$ induces a canonical inclusion $\iota:G(Y)\to G(X)$. Actually, one can easily check the identification $G(Y)=G(X)|^{\beta Y}_{\beta Y}$.
@@
Write $V_{Y}$ the $G(X)$-invariant open subset in $\beta X$. Notice that $\mathcal{I}(X,V_{Y})$ is isomorphic to the ideal of $C^{*}(X)$ generated by $Y$, denoted by $C^{*}(X,Y)$ as in Higson et al. [1993]. Thus, one has that
@@
With notation as above, the canonical inclusion map $\iota:C^{*}(Y)\cap\mathcal{I}\to\mathcal{I}(X,V_{Y})\cap\mathcal{I}$ induces an isomorphism in $K$-theory
@@
By 2.9, it induces an isomorphism on $K$-theory, i.e.,
@@
Let $X$ be a metric space with bounded geometry. Fix a $G(X)$-invariant open subset $U\subseteq\beta X$. For any subspace $Y\subseteq X$, denote $U^{\beta Y}=\beta Y\cap U$. If the inclusion $i:\mathcal{I}(Y,U^{\beta Y})\to\mathcal{G}(Y,U^{\beta Y})$ induces an isomorphism in $K$-theory for any sparse subspaces $Y\subseteq X$, then the inclusion $i:\mathcal{I}(X,U)\to\mathcal{G}(X,U)$ induces an isomorphism in $K$-theory, i.e.,
$i_{*}:K_{*}(\mathcal{I}(X,U))\to K_{*}(\mathcal{G}(X,U))$
is an isomorphism.
@@
Take $\{Y,Z\}$ to be an $\omega$-excisive cover of $X$ such that $Y,Z$ and $Y\cap Z$ are all sparse. Since $Y\cup Z=X$, we have $V_{Y}\cup V_{Z}=\beta X$ by the definition of the Stone-Čech compactification, where $V_{Y}$ and $V_{Z}$ are $G(X)$-invariant open subset generated by $Y$ and $Z$, respectively. It is proved in Higson et al. [1993] that
@@
induces an isomorphism on $K$-theory for $W=Y,Z$ and $Y\cap Z$ which are all sparse. By 2.11, we conclude that $i_{*}$ is identified with
@@
By 2.10, the map $i_{*}$ is further identified with
@@
Therefore, it suffices to prove $i_{*}$ is an isomorphism for all sparse subspaces of $X$, as the general case will then follow. This completes the proof of the proposition. ∎
@@ 3. The $\ell^{p}$-Bott-Dirac construction and twisted ghostly ideals
In the following subsections, we shall prove 2.6 in several steps. First of all, we shall briefly recall the definition of the Bott-Dirac operator in finite-dimensional $\ell^{p}$-spaces. Please refer to Wang et al. [2024] for a detailed introduction.
@@
For any $v\in(E,\|\cdot\|_{p})$, we define the Bott generator to be the unbounded function
@@
We summarize the facts we will need about the Bott-Drac operator in the next result. A detailed proof for the proposition can be found in [Wang et al., 2024, Lemma 2.7]:
Proposition 3.1 (Wang et al. [2024]).
@@
In Wang et al. [2024], the identity map $f(v)=v$ used in the construction of the Clifford operator in [Willett and Yu, 2020, Section 12.1] is replaced by the extended Mazur map. Consequently, the Bott-Dirac operator remains well-behaved. This is primarily due to the fact that $\Psi$ is a Lipschitz map, see [Wang et al., 2024, Lemma 2.6].
Lemma 3.2.
Assume $p\geq 2$. The extended Mazur map $\Psi:(E,\|\cdot\|_{2})\to(E,\|\cdot\|_{p})$ is Lipschitz, i.e.,
$\|\Psi(v)-\Psi(u)\|_{2}\leq L\|v-u\|_{p},$
for any $u,v\in(E,\|\cdot\|_{p})$, where $L=1+p\cdot 2^{p}$.
It is worth noting that 3.2 is crucial for the spectral analysis of $B_{s,v}$. This lemma is also essential for controlling the support of the bounded Bott-Dirac operator $F_{s,v}$, which we define below and will use frequently throughout the paper.
@@
Notice that $F_{s,v}$ is a bounded operator on $L^{2}(E,\mathfrak{h}_{E})$. There is a natural action of $C_{0}(E)$ on $L^{2}(E,\mathfrak{h}_{E})$, given by pointwise multiplication. Consequently, for any $p\in[1,\infty)$, we can discuss the propagation of Fs,vF{s,v} in the $E$-direction with respect to the norm $\|\cdot\|_{p}$. We shall denote this number by $\textup{Prop}_{E,p}(F_{s,v})$.
For any $v\in E$ and $R>0$, we denote $\chi_{v,R}:=\chi_{B(v,R)}$ for simplicity. The detailed properties of $F_{s,v}$ is summarized in [Willett and Yu, 2020, Proposition 12.1.10] for the case when $p=2$ and [Wang et al., 2024, Proposition 2.17] for the case when $p\geq 2$. Here we simplify this proposition as follows.
Proposition 3.4 (Willett and Yu [2020]; Wang et al. [2024]).
For each $\varepsilon>0$, there exists an odd function $\Phi:\mathbb{R}\to[-1,1]$ with $\lim_{t\to\pm\infty}\Phi(t)=\pm 1$, satisfying the following properties:
@@
    For all $s\in[1,\infty)$ and $v\in E$, $\|F_{s,v}-\Phi(B_{s,v})\|<\varepsilon$. Moreover, there exists $R_{0}>0$ such that $\textup{Prop}_{E,p}(\Phi(B_{s,v}))\leq s^{-1}R_{0}$ for all $s,v$ and $p\geq 2$.
@@
    For all $s\in[1,\infty)$, the operators $\Phi(B_{s,v})^{2}-1$ and $\Phi(B_{s,v})-\Phi(B_{s,w})$ are compact. Furthermore, the map $v\mapsto\Phi(B_{s,v})$ is Lipschitz continuous with respect to the $p$-norm on $E$, i.e., Φ(B_s,v)-Φ(B_s,w)∥≤cv-w_p.
@@
    The map $s\mapsto\Phi(B_{s,v})$ is strongly continuous. Moreover, for any $r>0$, the families {sΦ(B_s,v)^2-1}_vE  and {sΦ(B_s,v)-Φ(B_s,w)}_v-w_pr are norm equi-continuous on compact subsets of $1,\infty)$.
@@
    For any $r>0$, there exists $R_{\varepsilon,r}>0$ (independent of $\dim E$) such that for large $R\geq R_{\varepsilon,r}$, $v,w\in E$ with $\|v-w\|_{p}\leq r$ and $s\in[L\cdot(\operatorname{dim}E_{n})^{2},\infty)$, the operators satisfy the following norm estimates: (Φ(B_s,v)^2-1)(1-χ_v,R)¡ε and(Φ(B_s,v)-Φ(B_s,w))(1-χ_v,R)¡ε.
@@
The first step ensures that, without loss of generality, we may assume $X=\bigsqcup_{n=0}^{\infty}X_{n}$ is a sparse space that admits a coarse embedding into $\ell^{p}$. By [Willett and Yu, 2020, Lemma 12.5.4], we can assume the distance between $X_{n}$ and $X_{m}$ is infinite whenever $n\neq m$ (called separated disjoint union of $\{X_{n}\}_{n\in\mathbb{N}}$). In this case, the Hilbert space $\mathcal{H}_{X}$ has a canonical decomposition $\mathcal{H}_{X}=\bigoplus_{n\in\mathbb{N}}\mathcal{H}_{X_{n}}$. The Roe algebra $C^{*}(X)$ is therefore a subalgebra of $\prod_{n\in\mathbb{N}}C^{*}(X_{n})$
@@
Next, we will use the bounded Bott-Dirac map from 3.3 to construct the index map:
@@
For each $s\in[1,\infty)$, $F_{s,v,E_{n}}:L^{2}(E_{n},\mathfrak{h}_{E_{n}})\to L^{2}(E_{n},\mathfrak{h}_{E_{n}})$ is the bounded operator as in 3.3. Set $s_{n}:=L(\operatorname{dim}E_{n})^{2}$, where $L$ is the Lipschtiz constant in 3.2.
@@
Analogously, if $\Phi$ has the properties in 3.4 for some $\varepsilon>0$, we write $F^{\Phi}$, $F^{\Phi}_{s}$ and $F^{\Phi}_{s,n}$ for the operators built in the same way as $F$, $F_{s}$ and $F_{s,n}$, but replacing $F_{s,v,E_{n}}$ with $\Phi(B_{s,v,E_{n}})$.
By 3.4, the function $F$ is strong *-continuous with respect to $s\in[1,\infty)$. To introduce the index map, we also need the following lemma. The proof of the lemma can be found in [Wang et al., 2024, Proposition 2.19] and [Willett and Yu, 2020, Section 12.3]. We will omit some details here.
@@
It should be noted that in 3.6, the parameter $s$ is shifted to the right by $s_{n}$ (depending on the dimension) to ensure the validity of the preceding 3.7 applying (4) in 3.4.
By [Willett and Yu, 2020, Corollary 2.7.4], a $K_{0}$-theory element of $C^{*}(X)$ can be represented by a single projection $[P]$. For any projection $P\in C^{*}(\mathcal{H}_{X})$, $PF_{s}P$ is an odd, self-adjoint operator on the graded Hilbert space $L^{2}[1,\infty)\otimes\mathcal{H}_{X,E}$. We can build an index class of $K_{0}(PA(X,E)P)$ by using $PF_{s}P$ and 3.7. We can decompose $PF_{s}P$ as a matrix with respect to the grading on each $\mathcal{H}_{X,E}$ and $PF_{s}P$ has the form
@@
The index map is a well-defined group homomorphism. For each $s\in[1,\infty)$, define $\iota^{s}:A(X,E)\to C^{*}(X,H_{X,E})$ to be the evaluation map $(T_{s})\mapsto T_{s}$. The following proposition is proved in [Wang et al., 2024, Proposition 2.20], also see [Willett and Yu, 2020, Section 12.3]. We shall provide a detailed proof in 3.13.
@@
We will only prove the case for the map $s\mapsto T(F_{s}^{2}-1)$ by using (7) in 3.4, as the argument for $T,F_{s}]$ is analogous by using (4) in 3.4. By 3.7, we already have these two functions belong to $A(X,E)$.
@@
Combining this with 3.11, we have that